OpenStudy (anonymous):

what is principal when interest is $10, interest rate is 8%, and time if 2 years?

hartnn (hartnn):

Simple interest, right ? just use I = P*N*R/100 where I =10, N=2,R=8 find P from here

OpenStudy (anonymous):

how do I get the principal, what is the formula?

hartnn (hartnn):

I = P*N*R/100 P = Principal, which u need to find I= interest = 10 N= number of years = 2 R=rate f interest =8

OpenStudy (anonymous):

is it principal divided by rate x time?

hartnn (hartnn):

so, 10 = P*2*8/100 can u find P from here ?

OpenStudy (anonymous):

so is it $10 divided by .16 or $10 divided by 2.16, I'm still confused , sorry

hartnn (hartnn):

P = 10*100/(2*8) =...

OpenStudy (anonymous):

ok so P equals $62.50 right?

hartnn (hartnn):

yep, correct :)

OpenStudy (anonymous):

ok, thx

hartnn (hartnn):

welcome ^_^